OpenStudy (anonymous):

Solve for x: |x|-8 =-5

geerky42 (geerky42):

try to isolate |x| first should be easy, right?

OpenStudy (anonymous):

Idk I don't understand this

OpenStudy (agreene):

abs can be thought of as an operation that you can move. So if you isolate the variables from the numerical components you can abs the numeric value. \[\left| x \right|-8=-5\] \[\left| x \right| = -5+8\] \[\left| x \right| = 3\] \[x = \left| 3 \right|\] \[x=3\]

OpenStudy (anonymous):

But shouldn't there be another answer

OpenStudy (agreene):

well, if the answer on the right was negative when |x|=-5+8 then yes, there could be two answers (the negative and positive answer) however, this is more of a nameing/teaching thing and isnt actually true.